The charge for electricity is 3 cents per unit for lighting and 1/2 cents per unit for heating.A man's bill for a quarter should have been $8. by mistakes his lighting was charged at 1/2cents per unit and his heating at 3 cents per unit. the amount of the bi1l was $13. use any matrix method to find how many units of electricity he used altogether
Answer by stanbon(75887)

The charge for electricity is 3 cents per unit for lighting and 1/2 cents per unit for heating.
--------------------------
A man's bill for a quarter should have been $8.
-----------------------------------------------------
By mistakes his lighting was charged at 1/2 cents per unit and his heating at 3 cents per unit. the amount of the bi1l was $13.
-----------------------------------------------------
Use any matrix method to find how many units of electricity he used altogether
===============================================
Equations:
Proper charges: 3L + (1/2)H = 800 cents
Wrong Charges: (1/2)L+ 3H = 1300 cents
-----------------------------------------------
Multiply thru 1st by by 2.
Multiply thru 2nd by 12
----
6L + H = 1600
6L = 18H = 12*1300
----------------------------
Subtract and solve for "H":
17H = 14000
H = 823.53
----
Solve for L:
3L + (1/2)H = 800
3L + 411.76 = 800
3L = 388.24
L = 129.4``
----
Total # of units = 823.53 + 129.41 = 952.94
=============================================
Cheers,
Stan H.